# Phonon dynamics and chiral modes in the two-dimensional square-octagon lattice

### v1

@@ Abstract
Chiral phonons, originally identified in two-dimensional hexagonal lattices and later extended to kagome, square and other lattices, have been extensively studied as manifestations of broken inversion and time-reversal symmetries in vibrational dynamics. In this work, we investigate the vibrational dynamics of the two-dimensional square-octagon lattice using a spring-mass model with central-force interactions. The model incorporates mass contrast and variable coupling strengths among nearest, next-nearest, and third-nearest neighbors. From the dynamical matrix, we obtain the phonon dispersion relations and identify tunable phononic band gaps governed by both mass and spring-constant ratios. The angular dependence of phase and group velocities is analyzed to reveal the pronounced anisotropy inherent to this lattice geometry. We also examine the distinctive features of the square-octagon geometry, including flat-band anomalies in the density of states and anisotropic sound propagation induced by longer-range couplings. In addition, we explore the emergence of chiral phonons by introducing time reversal symmetry breaking term in the dynamical matrix and to elucidate their optical signatures, we construct a minimal model to study infrared circular dichroism arising from chiral phonon modes.
@@ I Introduction
Phonons, first introduced in the 1930s [1], are fundamental to understanding the thermal, mechanical, and transport properties of crystalline solids. The field of lattice dynamics, which investigates these quantized collective excitations, has since become an essential branch of condensed matter physics. Chiral phonons proposed by Zhang et.al [2] and later experimentally observed [3, 4, 5] has changed our understanding of how atomic vibrations can interact with electronic and magnetic degrees of freedom and thus has attracted great interest in recent years[6, 7, 8, 9, 10]. These quantized lattice vibrations with circular polarization and non-zero angular momentum are distinguished from conventional linearly polarized phonons by their rotational character. The emergence of chiral phonons requires primarily the breaking of time-reversal and/or inversion symmetries. When these symmetries are broken simultaneously, truly chiral phonons can emerge with well-defined handedness[11, 12, 13, 14].
@@
In this work we investigate the phonon properties of the square-octagon lattice within a classical spring-mass framework. This lattice, comprising of four atoms per unit cell and known for hosting multiple flat electronic bands and van Hove singularities (vHS) [39], remains largely unexplored from the perspective of lattice vibrations. We calculate the phonon dispersion for various mass ratios and spring constants by incorporating interactions beyond the nearest neighbors, thereby capturing the influence of longer-range couplings on the vibrational spectrum, along with the corresponding density of states (DOS), mode profiles, and the angular dependence of both phase and group velocities. Furthermore, we examine the emergence of chiral phonons in this lattice when time-reversal symmetry ($\mathcal{T}$) is explicitly broken. To highlight the connection between phonon helicity and optical activity we also construct a minimal model to study infrared circular dichroism arising from chiral phonon modes.
@@ VI Code and data availability
The relevant codes used to generate data for the current study are available from the corresponding author upon reasonable request.

### v2

@@ Abstract
Chiral phonons, originally identified in two-dimensional hexagonal lattices and later extended to kagome, square, and other lattices, have been extensively studied as manifestations of broken inversion and time-reversal symmetries in vibrational dynamics. In this work, we investigate the vibrational dynamics of the two-dimensional square-octagon lattice using a spring-mass model with central-force interactions. The model incorporates mass contrast and variable coupling strengths among nearest, next-nearest, and third-nearest neighbors. From the dynamical matrix, we obtain the phonon dispersion relations and identify tunable phononic band gaps governed by both mass and spring-constant ratios. The angular dependence of phase and group velocities is analyzed to reveal the pronounced anisotropy inherent to this lattice geometry. We also examine the distinctive features of the square-octagon geometry, including flat-band anomalies in the density of states and anisotropic sound propagation induced by longer-range couplings. In addition, we explore the emergence of chiral phonons by introducing a time reversal symmetry-breaking term in the dynamical matrix, and to elucidate their optical signatures, we construct a minimal model to study infrared circular dichroism arising from chiral phonon modes.
@@ I Introduction
Phonons, first introduced in the 1930s [1], are fundamental to understanding the thermal, mechanical, and transport properties of crystalline solids. The field of lattice dynamics, which investigates these quantized collective excitations, has since become an essential branch of condensed matter physics. Chiral phonons proposed by Zhang et.al [2] and later experimentally observed [3, 4, 5] has changed our understanding of how atomic vibrations can interact with electronic and magnetic degrees of freedom and thus has attracted great interest in recent years[6, 7, 8, 9, 10, 11, 12]. These quantized lattice vibrations with circular polarization and non-zero angular momentum are distinguished from conventional linearly polarized phonons by their rotational character. The emergence of chiral phonons requires primarily the breaking of time-reversal and/or inversion symmetries. When these symmetries are broken simultaneously, truly chiral phonons can emerge with well-defined handedness[13, 14, 15, 16].
@@
In this work we investigate the phonon properties of the square-octagon lattice within a classical spring-mass framework. This lattice, comprising of four atoms per unit cell and known for hosting multiple flat electronic bands and van Hove singularities (vHS) [41, 42], remains largely unexplored from the perspective of lattice vibrations. We calculate the phonon dispersion for various mass ratios and spring constants by incorporating interactions beyond the nearest neighbors, thereby capturing the influence of longer-range couplings on the vibrational spectrum, along with the corresponding density of states (DOS), mode profiles, and the angular dependence of both phase and group velocities. Furthermore, we examine the emergence of chiral phonons in this lattice when time-reversal symmetry ($\mathcal{T}$) is explicitly broken. A recent study [43] has done group theory analysis offering experimental benchmarks using circularly polarized Raman scattering for identifying crystalline chirality. To highlight the connection between phonon helicity and optical activity we also construct a minimal model to study infrared circular dichroism arising from chiral phonon modes.
@@ VI Code and data availability
The relevant codes used to generate data for the current study are available at Zenodo [49].